# FCS-HGNN: Flexible Multi-type Community Search in Heterogeneous Information Networks

### v2

@@ 4.5. Training and Online Query
4.5.1. Training
For the training stage, following previous work (Gao et al., 2021; Jiang et al., 2022b), we define the community search problem as a binary node classification problem, where the goal is to train the model to determine whether each node belongs to the community under query $q$. We denote $y\in\{0,1\}$ as the true label of each node under query $q$. Specifically, $y_{i}=1$ if and only if node $v_{i}$ belongs to community $\mathcal{C}_{q}$; otherwise, $y_{i}=0$. Then, we utilize the Binary Cross Entropy (BCE) function as the loss function to minimize the discrepancy between the predicted and ground-truth labels:
where $y_{i}$ and $p_{i}$ are the ground-truth label and predicted probability of node $v_{i}$ under query $q$, respectively. The label of $y_{i}$ will change with different queries $q$. With the training data $\mathcal{D}$, we update the parameters of FCS-HGNN according to the Equation 9.
Output: The optimal parameters of FCS-HGNN
1Initialize the parameters of FCS-HGNN;
2 while not converge do
3       for $(q,pos(q),neg(q))\in\mathcal{D}$ do
             /* Heterogeneous & Query Information Preprocessing */
4             $\mathcal{H}_{h}\leftarrow$ type-shared feature projection by Eq. 1;
5             $\mathcal{H}_{q}\leftarrow$ query-dependent information by Eq. 2;
             /* Information Integration */
9             $l=1$;
10             while $l<L$ do
14                   $l=l+1$;
15
16            $p\leftarrow$ calculated probability of each node by Eq. 8;
17             Update the parameters according to the loss in Eq. 9;
18
19
return The optimal parameters of FCS-HGNN.
Algorithm 2 The training process of FCS-HGNN
Algorithm 2 outlines the training process of our proposed FCS-HGNN. In the loop, we first deal with the heterogeneity of nodes through type-shared feature projection to obtain $\mathcal{H}_{h}$ and construct the query-dependent graph $\mathcal{H}_{q}$ based on the query $q$ (lines 4-5). Then, we extract complementary information from different views through the heterogeneous encoder and query encoder, and integrate them via the feature fusion module (lines 6-14). Finally, we compute the probability of each node belonging to the community and update the model parameters (Lines 15-16). It is noteworthy that $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$ differ solely in terms of node features, sharing the same graph structure. This characteristic ensures that there is no additional computational cost overhead incurred.
Furthermore, in our proposed FCS-HGNN, the single-type community is considered as a special case of the multi-type community, where $|\mathcal{S}_{\mathcal{A}}|=1$. During the training phase, the node types of the target community are agnostic to the model. This enables the model to focus on learning the inherent features of the HIN, query-dependent information, and the community patterns present in the training data. Then, we leverage the predicted probability of each node belonging to the community and the search algorithm to focus on the user-specified node types $\mathcal{S}_{\mathcal{A}}$. In other words, the training phase focuses on learning community patterns and the search algorithm considers user-specified node types. In this way, we are able to flexibly handle both single-type and multi-type communities, as shown in Table 1. Furthermore, our proposed FCS-HGNN significantly alleviates the burden of predefined metapaths and designing a series of relationship constraints.
4.5.2. Online Query
For the online query phase, we utilize the best model obtained during the training phase to execute queries instead of retraining the model for each query. For a new query $q_{new}$, we first employ the model to calculate the probability of each node belonging to the community. Then, we output the target community based on the search algorithm, as shown in Algorithm 1.
5. LARGE-SCALE FCS-HGNN
When dealing with large-scale graphs, traditional community search methods (Jiang et al., 2022a; Liu et al., 2023) usually improve efficiency by designing appropriate indexes. However, machine learning (ML)-based community search methods (Gao et al., 2021; Jiang et al., 2022b) face significant challenges when handling large-scale graphs due to the substantial computational costs. To address this issue, we introduce the neighbor sampling algorithm to enhance training efficiency and propose the depth-based heuristic strategy to improve query efficiency, resulting in the optimized algorithm named LS-FCS-HGNN.
5.1. Neighbor Sampling
@@
To improve the training efficiency, we introduce a neighbor sampling algorithm, which recursively samples a fixed number of neighbors for each node to form a subgraph for training (Hamilton et al., 2017). Algorithm 3 outlines the neighbor sampling algorithm, where $S_{l}$ represents the set of nodes sampled in the $l$-th layer, $N_{v_{i}}^{l-1}$ represents the set of nodes sampled from the neighbors of $v_{i}$, utilized for training in the $(l-1)$-th layer, and fanouts $\{f_{1},f_{2},...,f_{L}\}$ are hyperparameters indicating the number of neighbors that should be sampled at each layer. We initialize $S_{L}$ with the labeled nodes, as the probabilities of these nodes are ultimately used to compute the loss in Equation 9 (Line 1). Then, for each node $v_{i}$ in $S_{l}$, we sample a fixed number $f_{l-1}$ of nodes from the neighbors of $v_{i}$, and add them to $S_{l-1}$ (Lines 2-6). Finally, we extract the corresponding subgraph $\mathcal{H}^{\prime}$ from $\mathcal{H}$ based on the nodes sampled at each layer (Line 7).
@@ 4             $S_{l-1}\leftarrow S_{l-1}\cup\{N_{v_{i}}^{l-1}\}$;
Input: Training data $\mathcal{D}$, HIN $\mathcal{H}$, the number of multi-head attention $K$, the number of layer $L$, Fanouts $f_{1},f_{2},...,f_{L}$
Output: The optimal parameters of LS-FCS-HGNN
1Initialize the parameters of LS-FCS-HGNN;
@@
             /* Neighbor Sampling */
@@
             /* Training model */
5             Line 4-16 in Algorithm 2 with $\mathcal{H^{\prime}}$;
6
7
return The optimal parameters for LS-FCS-HGNN.
Algorithm 4 The training process of LS-FCS-HGNN
In addition, based on the neighbor sampling (Algorithm 3), we outline the training process of LS-FCS-HGNN in Algorithm 4. Compared to Algorithm 2, LS-FCS-HGNN incorporates the process of neighbor sampling (Line 4) and utilizes $\mathcal{H}^{\prime}$ instead of $\mathcal{H}$ for subsequent training (Line 5). It is worth noting that we perform neighbor sampling at each epoch to obtain a variety of subgraphs, which allows the model to observe various structures and thereby comprehensively understand the entire community pattern. Therefore, the neighbor sampling strategy effectively reduces the computational costs of community search in large-scale graphs while maintaining the community effectiveness. Subsequent experiments in Section 6.2.3 further demonstrate the effectiveness and efficiency of the neighbor sampling strategy.
5.2. Depth-based Heuristic Strategy
To improve the query efficiency in the large-scale graphs, we propose the depth-based heuristic search strategy. The core hypothesis of our proposed depth-based heuristic strategy is that: the nodes closer to the query node are more likely to belong to the same community. This assumption is grounded in real-life observations. In many cases, individuals within a community tend to have stronger connections with each other compared to individuals outside the community. In other words, nodes within the same community typically have more edges and shorter paths between them. Therefore, we propose the depth-based heuristic strategy to limit the exploration depth, reducing the search on distant nodes and thereby enhancing search efficiency.
Algorithm 5 outlines the depth-based heuristic search algorithm. Compared to Algorithm 1, the queue $\mathcal{Q}$ not only stores node information but also records the depth of each node from the query node $q$ (Line 3). Since the depth explored by BFS is always the closest distance from the node to the query node, we use the depth to determine whether to explore the node (Line 6). Then, we do not restrict the enqueue conditions and add a node to the target community if the node satisfies the community conditions (Lines 9-11).
Output: The target community $\mathcal{C}_{q}$
$\mathcal{V}\leftarrow\varnothing$ // recorded visited nodes
$\mathcal{Q}\leftarrow(q,0)$ // Add $q$ and depth 0 into the queue
2 while $\mathcal{Q}$ is not empty do
3       (node $v_{i}$, depth $d$) $\leftarrow$ dequeue from $\mathcal{Q}$;
4       if $v_{i}\notin\mathcal{V}$ and $d<d_{max}$ then
             $\mathcal{V}\leftarrow\mathcal{V}\cup\{v_{i}\}$ // add visited nodes
5             foreach $v_{j}\in\mathcal{N}_{i}$ do
                   $\mathcal{Q}\leftarrow(v_{j},d+1)$ // add new node for BFS
6                   if $\psi(v_{j})\in\mathcal{S}_{\mathcal{A}}$ and $p_{j}>\gamma$ then
7                         $\mathcal{C}_{q}\leftarrow\mathcal{C}_{q}\cup\{v_{j}\}$;
8
9
10
11
return The target community $\mathcal{C}_{q}$.
Algorithm 5 Depth-based heuristic search algorithm
@@ 5.3. Time Complexity
In the following, we analyze the time complexity of our proposed algorithms to analyse the efficiency of FCS-HGNN and LS-FCS-HGNN.
Theorem 1 ().
Let $T$ denote the number of training iterations, $|V|$ represent the number of nodes, $|E|$ represent the number of edges, $|\mathcal{D}|$ be the quantity of the training set, and $F$ and $F^{\prime}$ represent the feature dimensions before and after linear transformation, respectively. The time complexity for a layer of heterogeneous encoder or query encoder is $\mathcal{O}(|V|FF^{\prime}+|E|F^{\prime})$, where $|V|FF^{\prime}$ is the time complexity for the transformation of feature vectors for all nodes, and $|E|F^{\prime}$ is the time complexity for computing attention. For the FCS-HGNN with $L$ layers, the time complexity for each iteration would be $\mathcal{O}(L(|V|FF^{\prime}+|E|F^{\prime}))$. Thus, the time complexity of the training process of FCS-HGNN (Algorithm 2) is $\mathcal{O}(TL|\mathcal{D}|(|V|FF^{\prime}+|E|F^{\prime}))$, which primarily depends on the size of the graph, such as $|V|$ and $|E|$.
Theorem 2 ().
Let $|V|$ and $|E|$ represent the number of nodes and edges respectively, the time complexity of search algorithm (Algorithm 1) is $\mathcal{O}(|V|+|E|)$. Since we only need to perform a forward propagation during the inference phase, the time complexity of the query process for FCS-HGNN is $\mathcal{O}(L(|V|FF^{\prime}+|E|F^{\prime})+|V|+|E|)$.
Theorem 3 ().
Let LS-FCS-HGNN have $L$ layers and there are $n_{l}$ nodes to be sampled in the $l$-th layer. The time complexity of neighbor sampling (Algorithm 3) is $\mathcal{O}(L\sum_{l=1}^{L}n_{l})$, which primarily depends on the number of operations ($n_{l}$) performed in each layer.
Theorem 4 ().
Let $n$ and $e$ represent the total number of nodes and edges in the subgraph respectively, where $n=\sum_{l=1}^{L}n_{l}$. According to the theorem 1 and 3, the time complexity of the training process of LS-FCS-HGNN (Algorithm 4) is $\mathcal{O}(TL|\mathcal{D}|(nFF^{\prime}+eF^{\prime}+Ln))$, where $Ln$ is the time complexity of neighbor sampling.
Theorem 5 ().
The time complexity of depth-based heuristic search algorithm (Algorithm 5) is bounded by $\mathcal{O}(|V|+|E|)$, which depends on the maximum exploration depth $d_{max}$ as well as the graph structure. Thus, the time complexity of the query process of LS-FCS-HGNN is bounded by $\mathcal{O}(L(nFF^{\prime}+eF^{\prime}+Ln)+|V|+|E|)$.
@@ 6. EXPERIMENTS
In this section, we first elaborate on the experimental setup. Subsequently, we compare the effectiveness and the efficiency of our proposed methods against state-of-the-art methods in both single-type and multi-type community searches. Finally, we conduct extensive ablation studies and parameter sensitivity analyses on our proposed methods.
@@ 6.1. Experimental Setup
To comprehensively evaluate the performance of our proposed method, we conducted experiments on five real-world heterogeneous information networks, i.e., IMDB (Luo et al., 2021), DBLP (Fu et al., 2020), ACM (Luo et al., 2021), Freebase (Lv et al., 2021), and OGB-MAG (Hu et al., 2020). Table 2 reports the statistics of the datasets, where $|\mathcal{A}|$ and $|\mathcal{E}|$ denote the number of node types and edge types, respectively. Node features are necessary for all datasets. Notably, Freebase possesses a wealth of edge relationships, while OGB-MAG is a large-scale HIN derived from the OGB benchmark (Hu et al., 2020).
@@
In reality, it is challenging to manually collect labeled communities for training, especially multi-type communities in large-scale graphs. Thus, we adopt traditional community search methods in HINs to automatically label the single-type (Jiang et al., 2022a) or multi-type communities (Jian et al., 2020) as the ground truth. Specifically, the objective of (Jiang et al., 2022a) is to identify $(k,\Psi)$-NMC communities containing nodes of a single type, while (Jian et al., 2020) aims to search for relational communities containing nodes of different types. Since other traditional community search methods (Fang et al., 2020; Zhou et al., 2023) in HINs focus on different community patterns that are not consistent with those of (Jian et al., 2020) and (Jiang et al., 2022a), it would be unfair to compare these methods. However, machine learning (ML)-based methods do not rely on predefined community patterns; instead, they learn community patterns in a data-driven manner. Therefore, although these communities (Jian et al., 2020; Jiang et al., 2022a) impose strict constraints on community patterns, they provide a relatively fair comparison for ML-based methods in the absence of labeled communities in HINs. For these reasons, in this paper, we mainly compare with state-of-the-art ML-based methods. We randomly select one node from the groundtruth community as a query node $q$, randomly select three nodes as $pos(q)$, and three nodes outside of the groundtruth community as $neg(q)$. In order to evaluate the effectiveness of the algorithm for unseen queries, we partition the training, validation, and test sets with a ratio of 1:1:3, which minimizes the labeling information. For the single-type community, we designate one type of node as the primary node. For the DBLP dataset, we consider either author or paper nodes as the primary nodes, resulting in two variants: DBLP-A and DBLP-P, respectively. For the other datasets, we designate the primary node for single-type community according to (Luo et al., 2021). For the multi-type community, we set the paper and author nodes as primary nodes for DBLP, ACM, and OGB-MAG, movie and actor nodes as primary nodes for IMDB, film and people nodes as primary nodes for Freebase.
@@
To evaluate the quality of the identified communities, we employed three metrics: F1-score (Fang et al., 2023), Jaccard Similarity (Zhang et al., 2020), and Normalized Mutual Information (NMI) (Danon et al., 2005). These metrics are designed to assess different aspects of the identified communities (Li et al., 2023), aiding in comprehensive evaluation of various methods.
@@
For our model, we construct $L=2$ layers with dimensions of 64 or 128, and used Adam (Kingma and Ba, 2014) for optimization. As for other hyperparameters, we set the learning rate to 0.001, the regularization parameter to 1e-4, the dropout rate to 0.5, the number of attention heads $K$ to 8, and the fanouts to $f_{1}=20,f_{2}=10$. The threshold $\gamma$ is obtained by searching for the best F1 score from the validation set. For LS-FCS-HGNN, we traverse the maximum depth $d_{max}$ from 2 to 10 and record the best results. For all the baselines, we fine-tuned their parameters according to their original papers and recorded their best performance. All experiments were conducted on Ubuntu 20.04 equipped with 2.40 GHz Intel(R) Xeon(R) Gold 6240R CPUs, 512GB memory, and 4 NVIDIA Tesla V100 GPUs.
@@ 6.2. Community Search Performance
In this section, we conducted a comprehensive comparison with state-of-the-art methods for both single-type and multi-type community searches in HINs. Then, we compare the training and query efficiency with state-of-the-art methods and conduct a case study.
@@
Since the single-type community is a special case of the multi-type community, we first compare the effectiveness of our proposed algorithms in the context of the single-type community search. Table 3 displays the experimental results of single-type community search in HINs. The best results from each dataset are highlighted in bold, while the best results from the baseline method are underlined for comparison. We also indicate the case of ”out of memory” (o.o.m).
From Table 3, it is evident that our proposed methods outperform other state-of-the-art methods across all datasets, achieving average improvements of 11.08%/13.08%/18.68% on the F1/JS/NMI metrics, respectively. This demonstrates the effectiveness of our methods when identifying single-type community search in HINs. We observe that these baselines generally perform poorly on the NMI metric, indicating that they fail to capture the structure and node distribution of the target communities adequately. However, our method not only significantly surpasses the baselines on the NMI metric but also does not exhibit lopsided performance across the three metrics. This indicates that our proposed methods effectively learn the complex community structure in HINs, which is crucial for community search. Specifically, our methods can extract valuable information from two different views $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$ and dynamically integrate information through the feature fusion module. In each encoder, our proposed edge semantic attention mechanism effectively helps the model learn the complex community structure in the HIN during the information integration. Furthermore, we perform a comparative analysis between FCS-HGNN and LS-FCS-HGNN. It is illuminating to observe that LS-FCS-HGNN produced superior results in most datasets except for DBLP-A, demonstrating the effectiveness of the neighbor sampling and depth-based heuristic strategies. Under the guidance of the depth-based heuristic strategy, LS-FCS-HGNN eliminates the interference from distant nodes, achieving superior results compared to FCS-HGNN which employs a full-graph search. Simultaneously, this validates the reasonableness of our hypothesis that nodes closer to the query node are more likely to belong to the same community.
@@
As shown in Table 4, our proposed methods exhibit outstanding performance across all datasets, achieving average improvements of 11.21%/10.24%/11.81% on the F1/JS/NMI metrics, respectively. On one hand, the type-shared feature projection adeptly addresses the heterogeneity of nodes in the absence of meta-paths or relational constraints, which maps nodes from different feature spaces into a unified feature space, thereby enhancing information integration. On the other hand, the edge semantic attention mechanism selectively aggregates information from neighbors based on different relations, effectively leveraging fine-grained heterogeneous information and facilitating the model’s understanding of HINs. Our subsequent ablation experiments further confirmed this observation. Moreover, we are able to obtain similar conclusions to single-type community search: our proposed neighbor sampling (Algorithm 3) and depth-based heuristic search strategies (Algorithm 5) not only empower LS-FCS-HGNN to efficiently identify communities in large-scale graphs, but also improve the quality and precision of community identification. This further demonstrates the effectiveness of our methods.
@@
Figure 3 illustrates the training time per epoch for various methods on all datasets. As ICS-GNN requires retraining for each query, we did not record its result of training efficiency. On the OGB-MAG dataset, QD-GNN, COCLE, and FCS-HGNN run out of GPU memory. As shown in Figure 3, we observe that when the graph size is small, the training efficiency of each method is similar, but as the size increases, the differences in training efficiency become more apparent. In general, FCS-HGNN exhibits better training efficiency than COCLE and COCLEP. However, our proposed LS-FCS-HGNN demonstrates exceptional training efficiency, surpassing all competing methods. Especially as the graph size increases, the advantage of LS-FCS-HGNN’s training efficiency becomes increasingly significant.
@@
In addition to training efficiency, we also thoroughly compared the query efficiency of each method, as shown in Figure 4. It is worth noting that each grid on the y-axis of Figure 4 differs by a factor of 10. Since ICS-GNN retrains the whole model for each new query, it exhibits the lowest query efficiency. Similar to training efficiency, the query efficiency of FCS-HGNN significantly outperforms that of COCLE and COCLEP. However, with the assistance of our proposed depth-based heuristic search strategy, the query efficiency of LS-FCS-HGNN is significantly superior to all other methods. As the graph size continues to increase, this advantage becomes even more pronounced. This further demonstrates the effectiveness of our proposed depth-based heuristic search strategy: reducing exploration of distant nodes to enhance query efficiency while maintaining or even improving query accuracy (as shown in Table 3 and 4).
Figure 3. Training Efficiency.
Figure 4. Query Efficiency.
6.2.5. Case Study
We conducted a case study on the DBLP dataset to compare with traditional community search methods (Jiang et al., 2022a; Jian et al., 2020) and COCLEP (Li et al., 2023). In Figure 5, figures (a-c) depict cases of single-type communities, while (d-e) represent cases of multi-type communities. specifically, Figure 4(a) is a case of ground truth in single-type community by (Jian et al., 2020) with $k=3$ and $\Psi=\{(APTPA)\}$. Figure 4(d) is a case of ground truth in multi-type community by (Jiang et al., 2022a) with relational constraint $\{\text{author},\text{paper},k=3\}$. In Figure 5, the query node is marked in white color, the black nodes are the nodes in ground-truth community, and the red nodes represent nodes returned by other methods but not present in the ground-truth communities. In nodes of different shapes, circles represent authors, while squares represent papers. We observe that our method returns a community that is the closest to the ground-truth community. Compared to COCLEP, our method demonstrates a better understanding of the complex relationships and community information in HINs.
Refer to caption (a) groundtruth (Jiang et al., 2022a) Refer to caption (b) COCLEP Refer to caption (c) LS-FCS-HGNN Refer to caption (d) groundtruth (Jian et al., 2020) Refer to caption (e) COCLEP Refer to caption (f) LS-FCS-HGNN
Figure 5. Case Study.
6.3. Ablation Study and Parameter Sensitivity Analysis
In this section, we conduct comprehensive ablation studies to assess the effectiveness of each module, and perform the sensitivity analyses on the key hyperparameters.

### v3

@@ 4.4. Training and Online Query
4.4.1. Neighbor Sampling
@@
To improve the training efficiency especially for large-scale graphs, we introduce a neighbor sampling algorithm, which recursively samples a fixed number of neighbors for each node to form a subgraph for training (Hamilton et al., 2017). Algorithm 2 outlines the neighbor sampling algorithm, where $S_{l}$ represents the set of nodes sampled in the $l$-th layer, $N_{v_{i}}^{l-1}$ represents the set of nodes sampled from the neighbors of $v_{i}$, utilized for training in the $(l-1)$-th layer, and fanouts $\{f_{1},f_{2},...,f_{L}\}$ are hyperparameters indicating the number of neighbors that should be sampled at each layer. We initialize $S_{L}$ with the labeled nodes (Line 1), as the probabilities of these nodes are ultimately used to compute the loss in Equation 8. Then, for each node $v_{i}$ in $S_{l}$, we sample a fixed number $f_{l-1}$ of nodes from the neighbors of $v_{i}$, and add them to $S_{l-1}$ (Lines 2-6). Finally, we extract the corresponding subgraph $\mathcal{H}^{\prime}$ from $\mathcal{H}$ based on the nodes sampled at each layer (Line 7).
@@ 4             $S_{l-1}\leftarrow S_{l-1}\cup\{N_{v_{i}}^{l-1}\}$;
4.4.2. Training
For the training stage, following previous work (Gao et al., 2021; Jiang et al., 2022b), we define the community search problem as a binary node classification problem, where the goal is to train the model to determine whether each node belongs to the community under query $q$. We denote $y\in\{0,1\}$ as the true label of each node under query $q$. The label of $y_{i}$ will change with different queries $q$. Then, we utilize the Binary Cross Entropy (BCE) function as the loss function:
Output: The optimal parameters of FCS-HGNN
1Initialize the parameters of FCS-HGNN;
@@ 4.4.2. Training
             /* Special For LS-FCS-HGNN */
             // Neighbor Sampling
@@
             /* Common Part */
             // Heterogeneous & Query Information Preprocessing
5             $\mathcal{H}_{h}\leftarrow$ type-shared feature projection by Eq. 1;
6             $\mathcal{H}_{q}\leftarrow$ query-dependent information by Eq. 2;
             // Information Integration
10             $l=1$;
11             while $l<L$ do
15                   $l=l+1$;
16
17            $p\leftarrow$ calculated probability of each node by Eq. 7;
18             Update the parameters according to the loss in Eq. 8;
19
20
return The optimal parameters of (LS-)FCS-HGNN.
Algorithm 3 The training process of (LS-)FCS-HGNN
Algorithm 3 outlines the training process of both our proposed FCS-HGNN and LS-FCS-HGNN. It is noteworthy that $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$ differ solely in terms of node features, sharing the same graph structure, which ensuring that there is no additional computational overhead. Compared to FCS-HGNN, LS-FCS-HGNN first samples the subgraph $\mathcal{H}^{\prime}$ for training (Line 4). Furthermore, we perform neighbor sampling at each epoch to obtain a variety of subgraphs, which allows the model to observe various structures and thereby comprehensively understand the entire community pattern. In this way, the neighbor sampling strategy effectively reduces the computational overhead of community search in large-scale graphs while maintaining the community effectiveness. Subsequent experiments in Section 5.2 further demonstrate the effectiveness and efficiency of the neighbor sampling strategy.
4.4.3. Online Query
For the online query phase, we utilize the best model obtained during the training phase to execute queries, instead of retraining the model for each query (Gao et al., 2021). For a new query $q_{new}$, we first employ the model to calculate the probability of each node belonging to the community. Then, we output the target community based on the search algorithm, as shown in Algorithm 1.
@@ 4.5. Time Complexity
Theorem 1.
Let $T$ denote the number of training iterations, $|V|$ represent the number of nodes, $|E|$ represent the number of edges, $|\mathcal{D}|$ be the quantity of the training set, and $F$ and $F^{\prime}$ represent the feature dimensions before and after linear transformation, respectively. The time complexity for a layer of heterogeneous encoder or query encoder is $\mathcal{O}(|V|FF^{\prime}+|E|F^{\prime})$. For the FCS-HGNN with $L$ layers, the time complexity for each iteration would be $\mathcal{O}(L(|V|FF^{\prime}+|E|F^{\prime}))$. Thus, the time complexity of the training process of FCS-HGNN is $\mathcal{O}(TL|\mathcal{D}|(|V|FF^{\prime}+|E|F^{\prime}))$, which primarily depends on the size of the graph, such as $|V|$ and $|E|$, and the time complexity of the training process of LS-FCS-HGNN is $\mathcal{O}(TL|\mathcal{D}|(nFF^{\prime}+eF^{\prime}+Ln))$, where $Ln$ is the time complexity of neighbor sampling, $n=\sum_{l=1}^{L}n_{l}$.
Theorem 2.
The time complexity of depth-based heuristic search algorithm (Algorithm 1) is bounded by $\mathcal{O}(|V|+|E|)$, which depends on the maximum exploration depth $d_{max}$ as well as the graph structure. Thus, the time complexity of the query process of LS-FCS-HGNN is bounded by $\mathcal{O}(L(nFF^{\prime}+eF^{\prime}+Ln)+|V|+|E|)$, and the time complexity of the query process for FCS-HGNN is $\mathcal{O}(L(|V|FF^{\prime}+|E|F^{\prime})+|V|+|E|)$.
@@ 5.1. Experimental Setup
To comprehensively evaluate the performance of our proposed method, we conducted experiments on five real-world HINs, i.e., IMDB (Luo et al., 2021), DBLP (Fu et al., 2020), ACM (Luo et al., 2021), Freebase (Lv et al., 2021), and OGB-MAG (Hu et al., 2020). Table 2 reports the statistics of the datasets, where $|\mathcal{A}|$ and $|\mathcal{E}|$ denote the number of node types and edge types, respectively. Node features are necessary for all datasets. Notably, Freebase possesses a wealth of edge relationships, while OGB-MAG is a large-scale HIN derived from the OGB benchmark (Hu et al., 2020).
@@
In reality, it is challenging to manually collect labeled communities for training, especially multi-type communities in large-scale graphs. Thus, we adopt traditional community search methods in HINs to automatically label the single-type (Jiang et al., 2022a) or multi-type communities (Jian et al., 2020) as the ground truth. Specifically, the objective of (Jiang et al., 2022a) is to identify $(k,\Psi)$-NMC communities containing nodes of a single type, while (Jian et al., 2020) aims to search for relational communities containing nodes of different types. Since other traditional community search methods (Fang et al., 2020; Zhou et al., 2023) in HINs focus on different community patterns that are not consistent with those of (Jian et al., 2020) and (Jiang et al., 2022a), it would be unfair to compare these methods. ML-based methods do not rely on predefined community patterns. Instead, they learn community patterns in a data-driven manner. Therefore, although these communities (Jian et al., 2020; Jiang et al., 2022a) impose strict constraints on community patterns, they provide a relatively fair comparison for ML-based methods in the absence of labeled communities in HINs. For these reasons, we mainly compare with state-of-the-art ML-based methods. For the single-type community, we designate one type of node as the primary node. For the DBLP dataset, we consider either author or paper nodes as the primary nodes, resulting in two variants: DBLP-A and DBLP-P, respectively. For the other datasets, we designate the primary node for single-type community according to (Luo et al., 2021). For the multi-type community, we set the paper and author nodes as primary nodes for DBLP, ACM, and OGB-MAG, movie and actor nodes as primary nodes for IMDB, film and people nodes as primary nodes for Freebase.
@@
To evaluate the quality of the identified communities, we employed three metrics: F1-score (Fang et al., 2023), Jaccard Similarity (Zhang et al., 2020), and Normalized Mutual Information (NMI) (Danon et al., 2005).
@@
For our model, we construct $L=2$ layers with dimensions of 64 or 128, and used Adam (Kingma and Ba, 2014) for optimization. As for other hyperparameters, we set the learning rate to 0.001, the regularization parameter to 1e-4, the dropout rate to 0.5, the number of attention heads $K$ to 8, and the fanouts to $f_{1}=20,f_{2}=10$. The threshold $\gamma$ is obtained by searching for the best F1 score from the validation set. For LS-FCS-HGNN, we traverse the maximum depth $d_{max}$ from 2 to 10 and record the best results. For all the baselines, we fine-tuned their parameters according to their original papers and recorded their best performance.
@@ 5.2. Community Search Performance
Since the single-type community is a special case of the multi-type community, we first compare the effectiveness of our proposed algorithms in the context of the single-type community search. Table 3 displays the experimental results of single-type community search in HINs.
From Table 3, it is evident that our proposed methods outperform other state-of-the-art methods across all datasets, achieving average improvements of 11.08%/13.08%/18.68% on the F1/JS/NMI metrics, respectively. This demonstrates the effectiveness of our methods when identifying single-type community search in HINs. We observe that these baselines generally perform poorly on the NMI metric, indicating that they fail to capture the structure and node distribution of the target communities adequately. Our method not only significantly surpasses the baselines on the NMI metric but also does not exhibit lopsided performance across the three metrics. This indicates that our proposed methods effectively learn the complex community structure in HINs, which is crucial for community search. Furthermore, it is illuminating to observe that LS-FCS-HGNN produced superior results in most datasets except for DBLP-A, demonstrating the effectiveness of the neighbor sampling and depth-based heuristic strategies. Under the guidance of the depth-based heuristic strategy, LS-FCS-HGNN eliminates the interference from distant nodes, achieving superior results compared to FCS-HGNN which employs a full-graph search.
@@
As shown in Table 4, our proposed methods exhibit outstanding performance across all datasets, achieving average improvements of 11.21%/10.24%/11.81% on the F1/JS/NMI metrics, respectively. On one hand, the type-shared feature projection adeptly addresses the heterogeneity of nodes in the absence of meta-paths or relational constraints, which maps nodes from different feature spaces into a unified feature space, thereby enhancing information integration. On the other hand, the edge semantic attention mechanism selectively aggregates information from neighbors based on different relations, effectively leveraging fine-grained heterogeneous information and facilitating the model’s understanding of HINs. Our subsequent ablation experiments further confirmed this observation. Moreover, we are able to obtain similar conclusions to single-type community search: the neighbor sampling and depth-based heuristic search strategies not only empower LS-FCS-HGNN to efficiently identify communities in large-scale graphs, but also improve the quality and precision of community identification. This further demonstrates the effectiveness of our methods.
@@
Figure 3(a) illustrates the training time per epoch for various methods on all datasets. As ICS-GNN requires retraining for each query, we did not record its result of training efficiency. On the OGB-MAG dataset, QD-GNN, COCLE, and FCS-HGNN run out of GPU memory. As shown in Figure 3(a), we observe that when the graph size is small, the training efficiency of each method is similar. As the size increases, the differences in training efficiency become more apparent. In general, FCS-HGNN exhibits better training efficiency than COCLE and COCLEP. However, our proposed LS-FCS-HGNN demonstrates exceptional training efficiency, surpassing all competing methods. Especially as the graph size increases, the advantage of LS-FCS-HGNN’s training efficiency becomes increasingly significant.
@@
In addition to training efficiency, we also thoroughly compared the query efficiency of each method, as shown in Figure 3(b). It is worth noting that each grid on the y-axis of Figure 3(b) differs by a factor of 10. Since ICS-GNN retrains the whole model for each new query, it exhibits the lowest query efficiency. Similar to training efficiency, the query efficiency of FCS-HGNN significantly outperforms that of COCLE and COCLEP. However, with the assistance of the depth-based heuristic search strategy, the query efficiency of LS-FCS-HGNN is significantly superior to all other methods. As the graph size continues to increase, this advantage becomes even more pronounced. This further demonstrates the effectiveness of our proposed depth-based heuristic search strategy.
(a) Training Efficiency.
(b) Query Efficiency.
Figure 3. Efficiency comparison
5.3. Ablation Study
(a) DBLP
(b) ACM